Find the principal invested if the simple interest earned in 5 years by an investment at 5% per annum is $100
1
Expert's answer
2019-01-17T09:04:11-0500

If the simple interest earned in 5 years by an investment at 5% per annum is $100, then the principal invested is

I = Prt,

P = I/(rt) = 100/(0.05*5) = $400.